On May 19, 2008, Area-RAE and pDR instruments were calibrated and setup at 4 different site locations for indoor and perimeter air monitoring. All readings for Area-RAE monitoring, including H2S, HCN, VOCs, and LEL were below Action Levels. O2 levels were between 20.8% and 21%. pDR dust concentrations were below nuisance levels for all locations.
The crew completed the removal of PVC process piping from the Basement and Main Plating Room. Piping and floor debris were staged for placement into a hazardous waste roll-off. All site operations ceased at 1530 for the holiday break. Security was on site from 1530 May 19, 2008 - 0700 May 28, 2008.
START conducted daily inspections of outside waste containers from May 20-May 22, 2008. All outside waste containers were intact.
On May 28, 2008, Area-RAE and pDR instruments were calibrated and setup at 4 identified site locations for indoor and perimeter air monitoring. All readings for Area-RAE monitoring, including H2S, HCN, VOCs, and LEL were below Action Levels. O2 levels were between 20.8% and 21%. pDR dust concentrations were below nuisance levels for all locations.
A 20 cubic yard roll-off container was delivered for hazardous waste debris by K-Com Transportation of Fort Wayne, IN. Scrap steel vats, wood and floor scrapings were placed into the lined roll-off for disposal. A tanker truck, also supplied by K-Com Transportation, arrived on site and transferred 1,320 gallons of heavy metal contaminated liquids from the building. This was transported to EQ Detroit, a Treatment and Disposal Facility located in Belleville, MI. Security was on-site from 1730 May 28, 2008 through 0700 May 29, 2008.
On May 29, 2008, Area-RAE and pDR instruments were calibrated and setup at 2 identified site locations for indoor air monitoring. Outdoor air monitoring was not utilized due to moderate rain at the site. A Multi-RAE PID was utilized in the Exclusion Zone (EZ) for CO monitoring. All readings for Area-RAE monitoring, including H2S, HCN, VOCs, and LEL were below Action Levels. O2 levels were between 20.8% and 21%. CO levels were 0ppm. pDR dust concentrations were below nuisance levels for all locations. The crew completed loading the 20-cubic yard roll-off with hazardous waste from the building. Additional waste was prepared for arrival of a second hazardous waste roll-off box. Security was on site from 1530 May 29, 2008 through 0700 May 30, 2008.
On May 30, 2008, Area-RAE and pDR instruments were calibrated and setup at 2 identified site locations for indoor air monitoring. Outdoor air monitoring was not utilized due to light rain at the site.
A Multi-RAE PID was utilized in the Exclusion Zone (EZ) for CO monitoring. All readings for Area-RAE monitoring, including H2S, HCN, VOCs, and LEL were below Action Levels. O2 levels were between 20.8% and 21%. CO levels were 0ppm. pDR dust concentrations were below nuisance levels for all locations.
Empty vats continued to be cut apart for disposal into the hazardous waste roll-off container. Four 55 gallon drums of sludge were recovered from an empty vat identified as a nickel bath. The floor in the Main Plating Room was scraped as additional vats were removed from this room. The scrapings were placed into the hazardous waste roll-off container. Security was on site from 1530 May 29, 2008 through 0700 May 30, 2008.
|